# The Rate-Distortion-Perception-Classification Tradeoff: Joint Source Coding and Modulation via Inverse-Domain GANs

### v1

@@ IV-B Solving RDPC: Problem Formulation
    The source signal $\bm{X}\in\mathbb{R}^{n}$ is drawn from a two-class Gaussian mixture model, i.e.,
@@
    The encoder $e\,\mathrel{\mathop{\ordinarycolon}}\,\mathbb{R}^{n}\to\mathbb{R}^{m}$ and decoder $d\,\mathrel{\mathop{\ordinarycolon}}\,\mathbb{R}^{m}\to\mathbb{R}^{n}$ are linear, i.e., they are implemented by full-rank matrices $\bm{E}\in\mathbb{R}^{m\times n}$ and $\bm{D}\in\mathbb{R}^{n\times m}$, respectively. In particular, $p_{\bm{Y}|\bm{X}}$ and $p_{\widehat{X}|\widehat{Y}}$ in (2) are deterministic.
@@
    We use the mean-squared error (MSE) as a metric for distortion, i.e., $\Delta(\bm{X},\,\widehat{\bm{X}})=\|\bm{X}-\widehat{\bm{X}}\|_{2}^{2}$, and the Wasserstein-1 distance $W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})$ as a metric for perception, where $\bm{X}\sim p_{\bm{X}}$ and $\widehat{\bm{X}}\sim p_{\widehat{\bm{X}}}$.
@@
Note that a linear encoder and decoder as in 2) can model a linearization of a fully trained neural network around an optimal point; see, e.g., work on the neural tangent kernel [30, 31]. Assumptions 1) and 2) imply that the reconstructed signal $\widehat{\bm{X}}$ is also a Gaussian mixture model. To see this, first note that the output signal is $\widehat{\bm{Y}}=\bm{D}\big{(}\bm{E}\bm{X}+\bm{N}\big{)}$, where $\bm{N}\sim\mathcal{N}(\bm{0}_{m},\bm{\Sigma})$. Since the sum of two Gaussian random variables is also Gaussian, we obtain
@@
Note that the covariance matrix $\widehat{\bm{\Sigma}}\mathrel{\mathop{\ordinarycolon}}=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$ of both distributions in (IV-B) is rank deficient, since $\bm{D}\in\mathbb{R}^{n\times m}$ with $n>m$, making the distributions degenerate. Henceforth, $\widehat{\bm{\Sigma}}^{-1}$ will thus refer to the generalized inverse of $\widehat{\bm{\Sigma}}$. Specifically, let $\widehat{\bm{\Sigma}}=\bm{Q}\bm{\Lambda}\bm{Q}^{\top}$ be an eigenvalue decomposition of $\widehat{\bm{\Sigma}}$, with $\bm{\Lambda}=\text{Diag}(\lambda_{1},\ldots,\lambda_{n})$ being a diagonal matrix of eigenvalues. Define $\bm{\Lambda}^{-1}$ as the diagonal matrix with diagonal entries $1/\lambda_{i}$ if $\lambda_{i}>0$, and 00 otherwise. Then, $\widehat{\bm{\Sigma}}^{-1}\mathrel{\mathop{\ordinarycolon}}=\bm{Q}\bm{\Lambda}^{-1}\bm{Q}^{\top}$. Similarly, $|\widehat{\bm{\Sigma}}|$ is the product of the positive entries of $\bm{\Lambda}$.
Problem formulation. Under Assumption 2, (7) becomes
$R(D,P,C)=\begin{array}[t]{cl}\underset{\bm{E},\,\bm{\Sigma},\,\bm{D}}{\min}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\big{\|}\bm{X}-\widehat{\bm{X}}\big{\|}_{2}^{2}\big{]}\leq D\\[2.84544pt] &W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq P\\[2.84544pt] &\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\leq C\,,\end{array}$ (23)
where we omitted the dependence of $\widehat{\bm{X}}$ on $\bm{E}$ and $\bm{D}$ for simplicity. Despite the simplifications under Assumption 2, problem (23) is still challenging, and we will solve instead an approximation by relaxing its last two constraints. Before doing so, we analyze each constraint in detail.
@@
where $\widehat{\bm{\Sigma}}\mathrel{\mathop{\ordinarycolon}}=\bm{D}\big{(}\bm{E}\bm{E}^{\top}+\bm{\Sigma}\big{)}\bm{D}^{\top}$. Finally, the second term of the right-hand side of (25) can be rewritten for $l=0,1$ as
@@
Perception constraint. We now consider the perception constraint in (23), which upper bounds the Wasserstein-$1$ distance $W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})$ by $P$. Both $p_{\bm{X}}$ and $p_{\widehat{\bm{X}}}$ are Gaussian mixture models for which, to the best of our knowledge, there is no closed-form expression for their Wasserstein-$p$ distance. There is, however, a closed-form expression for the Wasserstein-$2$ distance between Gaussian distributions. Specifically, let $\bm{X}\sim p_{\bm{X}}=\mathcal{N}(\bm{\mu_{X}},\,\bm{\Sigma_{X}})$ and $\bm{Y}\sim p_{\bm{Y}}=\mathcal{N}(\bm{\mu_{Y}},\,\bm{\Sigma_{Y}})$ be two Gaussian random vectors with means $\bm{\mu_{X}},\,\bm{\mu_{Y}}\in\mathbb{R}^{n}$ and positive semidefinite covariance matrices $\bm{\Sigma_{X}},\,\bm{\Sigma_{Y}}\succeq\bm{0_{n\times n}}$. It can be shown that the squared Wasserstein-$2$ distance between them is [32, 33]
@@
Our objective is thus to upper bound $W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})$ as a function of $W_{2}(p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{0})$ and $W_{2}(p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{1})$, which we define as in footnote 4 [or (12)] with expected values conditioned on $H_{0}$ or $H_{1}$. We have the following result.
@@
Let $p_{\bm{X}}$ (resp. $p_{\widehat{\bm{X}}}$) be a Gaussian mixture model following (1) [resp. (IV-B)], in which the probability of hypothesis $H_{0}$ is $p_{0}$ and of hypothesis $H_{1}$ is $p_{1}=1-p_{0}$. Then,
@@
Classification constraint. We now address the last constraint of (23). As in assumption 2.4), we assume a Bayes classifier, which achieves a minimal probability of error. Such a probability, however, does not have a closed-form expression, but is upper bounded by the Bhattacharyya bound [34]. For a two-class Gaussian mixture model $\bm{X}\sim p_{0}\,\mathcal{N}(\bm{\mu}_{0},\,\bm{\Sigma}_{0})+p_{1}\,\mathcal{N}(\bm{\mu}_{1},\,\bm{\Sigma}_{1})$, the bound is
$\mathbb{P}\big{(}\text{error}^{\star}\big{)}\leq\sqrt{p_{0}\,p_{1}}\int_{\mathbb{R}^{n}}\sqrt{p_{\bm{X}|H_{0}}(\bm{x})\,p_{\bm{X}|H_{1}}(\bm{x})}\,\operatorname{d\!}\bm{x}$
@@
where ||$\cdot$|| ⋅ | is the determinant of a matrix, and $\text{error}^{\star}$ is the classification error achieved by a Bayes classifier. We apply (30) to $\bm{X}$ and $\widehat{\bm{X}}$, whose models are in (1) and (IV-B). Thus, $\bm{\mu}_{0}=\bm{0}_{n}$, $\bm{\mu}_{1}=\bm{D}\bm{E}\bm{c}_{n}$, and $\bm{\Sigma}_{0}=\bm{\Sigma}_{1}=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$. Hence,
$\epsilon_{c_{0}}(\bm{X},\,\widehat{\bm{X}})$ $=\mathbb{P}\big{(}\text{class}(\widehat{\bm{X}})\neq c_{0}\big{(}\widehat{\bm{X}}\big{)}\big{)}$
@@
So, in (23), rather than bounding $\epsilon_{c_{0}}(\bm{X},\,\widehat{\bm{X}})\leq C$, we impose instead that the right-hand side of (31) is upper bounded by $C$, which is equivalent to
@@
IV-C Solving RDPC: Heuristic Algorithm
@@
Design of $\widehat{\bm{\Sigma}}$. While the original signals in (1) have non-degenerate distributions, the decoded signals in (IV-B) have degenerate distributions. Specifically, assuming that $\bm{E}\in\mathbb{R}^{m\times n}$ and $\bm{D}\in\mathbb{R}^{n\times m}$ have full rank and that $\text{range}(\bm{E})\cap\text{null}(\bm{D})=\emptyset$, the output signals in (IV-B) live in an $m$-dimensional subspace. If the fixed vector $\bm{c}_{n}$, which represents the distance between $\bm{X}|H_{0}$ and $\bm{X}|H_{1}$, is orthogonal to that subspace (equivalently $\bm{D}\bm{E}\bm{c}_{n}=\bm{0}_{n}$), then $\widehat{\bm{X}}|H_{0}$ and $\widehat{\bm{X}}|H_{1}$ become indistinguishable. In this case, classification is impossible and perception is also undermined [note that the second term in (29) requires $\|\bm{D}\bm{E}\bm{c}_{n}-\bm{c}_{n}\|_{2}$ to be small].
To avoid this, we first generate the (degenerate) covariance matrix $\widehat{\bm{\Sigma}}\mathrel{\mathop{\ordinarycolon}}=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$ by guaranteeing that the distance between $\bm{X}|H_{0}$ and $\bm{X}|H_{1}$ is preserved after transmitting these signals through the channel. We achieve this by guaranteeing that $\bm{c}_{n}$ is an eigenvector of $\widehat{\bm{\Sigma}}$ associated to eigenvalue $1$, while the remaining eigenvectors are associated to eigenvalues of smaller magnitude. Specifically, we set $\widehat{\bm{\Sigma}}=\bm{Q}\bm{\Lambda}\bm{Q}^{\top}$, where the first column of $\bm{Q}$ is $\bm{c}_{n}$ and the remaining ones are the output of Gram-Schmidt orthogonalization. Also, $\bm{\Lambda}=\text{Diag}(1,\lambda_{2},\ldots,\lambda_{m},0,\ldots,0)$, with $\lambda_{i}$ being drawn uniformly at random from $[0,1]$, for $i=2,\ldots,m$.
@@
Finding $(\bm{E},\bm{D})$. With $\widehat{\bm{\Sigma}}$ fixed and assuming that, at iteration $k$, $\bm{\Sigma}=\bm{\Sigma_{k-1}}$ is also fixed, we seek a factorization $\widehat{\bm{\Sigma}}=\bm{D}\bm{E}\bm{E}^{\top}\bm{D}^{\top}+\bm{D}\bm{\Sigma_{k-1}}\bm{D^{\top}}$. We do so via an intuitive process that leads to a unique factorization. Specifically, we design $\bm{E}$ and $\bm{D}$ such that $\bm{D}\bm{E}\bm{E}^{\top}\bm{D}^{\top}$ is as close to the identity matrix as possible (to preserve signals passing through the channel), while $\bm{D}\bm{\Sigma_{k-1}}\bm{D}^{\top}$ is as small as possible (to mitigate the effects of noise). Also, we ensure the principal direction $\bm{c}_{n}$ is preserved: $\bm{D}\bm{E}\bm{c}_{n}\simeq\bm{c_{n}}$. These requirements, weighted equally, can be cast as a optimization problem:
@@
$\mathinner{\dfrac{\partial{}g(\bm{E},\bm{D})}{\partial{\bm{E}}}}$ $=\bm{D}^{\top}(\bm{D}\bm{E}\bm{c_{n}}-\bm{c_{n}})\bm{c_{n}}^{\top}$ (43a)
$\mathinner{\dfrac{\partial{}g(\bm{E},\bm{D})}{\partial{\bm{D}}}}$ $=4\bm{D}\bm{\Sigma_{k}}\bm{D}^{\top}\bm{D}\bm{\Sigma_{k-1}}+2(\bm{I_{n}}-\widehat{\bm{\Sigma}})\bm{D}\bm{\Sigma_{k-1}}$
@@
Finding $\bm{\Sigma}$. Once the encoder-decoder pair is fixed at $(\bm{E_{k}},\bm{D_{k}})$, we find the diagonal rate matrix $\bm{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\text{Diag}(\bm{\sigma})\mathrel{\mathop{\ordinarycolon}}=\text{Diag}(\sigma_{1},\ldots,\sigma_{m})$ by applying a barrier method [35] to (37), i.e., we solve a sequence of problems in $t$, each of which is
@@
where $\widehat{\bm{\Sigma}}_{\bm{k}}=\bm{D_{k}}\bm{E_{k}}\bm{E_{k}}^{\top}\bm{D_{k}}^{\top}+\bm{D_{k}}\bm{\Sigma}\bm{D_{k}}^{\top}$. In (45b), $D_{k}$ absorbs all the terms independent from $\bm{\Sigma}$ when we set $\mathbb{E}[\big{\|}\widehat{\bm{X}}-\bm{X}\big{\|}_{2}^{2}]\leq D$ in (27) [including $D$]. To obtain (45c), note that imposing the right-hand side of (29) to be smaller than $P$ is equivalent to $\big{\|}\widehat{\bm{\Sigma}}_{\bm{k}}^{\frac{1}{2}}-\bm{I_{n}}\big{\|}_{F}^{2}\leq\big{(}P-\|\bm{D_{k}}\bm{E_{k}}\bm{c_{n}}-\bm{c_{n}}\|_{2}\cdot p_{1}\big{)}^{2}=\mathrel{\mathop{\ordinarycolon}}P_{k}$. And $h_{p}(\bm{\Sigma})$ depends on $\bm{\Sigma}$ via $\widehat{\bm{\Sigma}}_{\bm{k}}$. Finally, (45d) is the direct application of the log-barrier function to (32).
To solve each instance of (44), we apply again gradient descent. While the gradients of $h_{r}$ in (45a) and $h_{d}$ in (45b) can be computed directly, namely $\operatorname{d\!}h_{r}(\bm{\sigma})/\operatorname{d\!}\sigma_{i}=-1/(\sigma_{i}^{2}+\sigma_{i})$, for $i=1,\ldots,m$, and $\nabla_{\bm{\sigma}}h_{d}(\bm{\Sigma})=-\text{diag}(\bm{D_{k}}^{\top}\bm{D_{k}})/[D_{k}-\text{tr}(\bm{D_{k}}\bm{\Sigma},\bm{D_{k}}^{\top})]$, where $\text{diag}(\cdot)$ extracts the diagonal entries of a matrix into a vector, computing the gradients of $h_{p}$ in (45c) and $h_{c}$ in (45d) is more laborious. Their expressions are shown in (IV-B) where, for simplicity, we omitted the iteration index.
@@
12:    Find $(\bm{E_{k}},\bm{D_{k}})$ via gradient descent applied to (42) [cf. (IV-C)]

### v2

@@ IV RDPC Tradeoff under GMM signals and linear encoder and decoder
    The source $\bm{X}\in\mathbb{R}^{n}$ is drawn from a two-class GMM:
@@
    The encoder $e\,:\,\mathbb{R}^{n}\to\mathbb{R}^{m}$ and decoder $d\,:\,\mathbb{R}^{m}\to\mathbb{R}^{n}$ are linear and deterministic, i.e., they are implemented by full-rank matrices $\bm{E}\in\mathbb{R}^{m\times n}$ and $\bm{D}\in\mathbb{R}^{n\times m}$.
@@
    We use the mean-squared error (MSE) as a metric for distortion, i.e., $\Delta(\bm{X},\,\widehat{\bm{X}})=\|\bm{X}-\widehat{\bm{X}}\|_{2}^{2}$, and the Wasserstein-1 distance [^3] The Wasserstein-$p$ distance between two probability measures $p_{\bm{X}},p_{\bm{Y}}$ in $\mathbb{R}^{n}$ is $W_{p}(p_{\bm{X}},p_{\bm{Y}})=\big{(}\underset{\gamma\in\Pi(p_{\bm{X}},p_{\bm{Y}})}{\inf}\mathbb{E}_{(\bm{X},\bm{Y})\sim\gamma}\big{[}\|\bm{X}-\bm{Y}\|_{2}^{p}\big{]}\big{)}^{1/p}$, where $\Pi\big{(}p_{\bm{X}},p_{\bm{Y}}\big{)}$ is the set of all joint distributions with marginals $p_{\bm{X}}$ and $p_{\bm{Y}}$, and $1\leq p\leq+\infty$. $W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})$ as a metric for perception, where $p_{\bm{X}}$ and $p_{\widehat{\bm{X}}}$ are the distributions of $\bm{X}$ and $\widehat{\bm{X}}$.
@@
Assumptions 1) and 2) imply that the reconstructed signal $\widehat{\bm{X}}$ is also a GMM. To see this, first note that the output signal is $\widehat{\bm{X}}=\bm{D}\big{(}\bm{E}\bm{X}+\bm{N}\big{)}$, where $\bm{N}\sim\mathcal{N}(\bm{0}_{m},\bm{\Sigma})$. Since the sum of two Gaussian random variables is also Gaussian, we obtain
@@
Note that $\bm{D}\in\mathbb{R}^{n\times m}$ has more rows than columns ($n>m$), making the covariance matrix $\widehat{\bm{\Sigma}}:=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$ in (13) rank-deficient, and thus both distributions in (13) degenerate. Henceforth, $\widehat{\bm{\Sigma}}^{-1}$ will thus refer to the generalized inverse of $\widehat{\bm{\Sigma}}$. Specifically, let $\widehat{\bm{\Sigma}}=\bm{Q}\bm{\Lambda}\bm{Q}^{\top}$ be an eigenvalue decomposition of $\widehat{\bm{\Sigma}}$, with $\bm{\Lambda}=\text{Diag}(\lambda_{1},\ldots,\lambda_{n})$ being a diagonal matrix of eigenvalues. Define $\bm{\Lambda}^{-1}$ as the diagonal matrix with diagonal entries $1/\lambda_{i}$ if $\lambda_{i}>0$, and 00 otherwise. Then, $\widehat{\bm{\Sigma}}^{-1}:=\bm{Q}\bm{\Lambda}^{-1}\bm{Q}^{\top}$. Similarly, the generalized determinant of $|\widehat{\bm{\Sigma}}|$ is the product of the positive entries of $\bm{\Lambda}$.
IV-A Problem formulation
Under Assumption 2, problem (7) becomes
$R(D,P,C)=\begin{array}[t]{cl}\underset{\bm{E},\bm{D},\bm{\Sigma}}{\min}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\big{\|}\bm{X}-\widehat{\bm{X}}\big{\|}_{2}^{2}\big{]}\leq D\\[2.84544pt] &W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq P\\[2.84544pt] &\mathbb{E}\Big{[}\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\Big{]}\leq C\,,\end{array}$ (18)
where we omitted the dependence of $\widehat{\bm{X}}$ on $\bm{E}$ and $\bm{D}$ for simplicity. Despite the simplifications made under Assumption 2, problem (18) is still challenging, and we will solve instead an approximation by relaxing its last two constraints. Before doing so, we analyze each constraint in detail.
@@ IV-A Problem formulation
where $\widehat{\bm{\Sigma}}:=\bm{D}\big{(}\bm{E}\bm{E}^{\top}+\bm{\Sigma}\big{)}\bm{D}^{\top}$. Finally, the second term of the right-hand side of (20) can be rewritten for $l=0,1$ as
@@
Perception constraint. We now consider the perception constraint in (18), which upper bounds the Wasserstein-$1$ distance $W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})$ by $P$. Both $p_{\bm{X}}$ and $p_{\widehat{\bm{X}}}$ are Gaussian mixture models for which, to the best of our knowledge, there is no closed-form expression for their Wasserstein-$p$ distance. There is, however, a closed-form expression for the Wasserstein-$2$ distance between Gaussian distributions. Specifically, let $\bm{X}\sim p_{\bm{X}}=\mathcal{N}(\bm{\mu_{X}},\,\bm{\Sigma_{X}})$ and $\bm{Y}\sim p_{\bm{Y}}=\mathcal{N}(\bm{\mu_{Y}},\,\bm{\Sigma_{Y}})$ be two Gaussian random vectors with means $\bm{\mu_{X}},\,\bm{\mu_{Y}}\in\mathbb{R}^{n}$ and positive semidefinite covariance matrices $\bm{\Sigma_{X}},\,\bm{\Sigma_{Y}}\succeq\bm{0}_{n\times n}$. It can be shown that the squared Wasserstein-$2$ distance between them is [Olkin82-distance, Dowson82-Frechet]
@@
Our objective is thus to upper bound $W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})$ as a function of $W_{2}(p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{0})$ and $W_{2}(p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{1})$, which we define as in footnote 3 [or, in dual form, as in (42) below] with expected values conditioned on $H_{0}$ or $H_{1}$. We have the following result.
@@
Let $p_{\bm{X}}$ (resp. $p_{\widehat{\bm{X}}}$) be a GMM modeled as (12) [resp. (13)], in which the probability of hypothesis $H_{0}$ is $p_{0}$ and of hypothesis $H_{1}$ is $p_{1}=1-p_{0}$. Then,
@@
Classification constraint. We now address the last constraint of (18). As in Assumption 2.4), we assume a Bayes classifier, which achieves a minimal probability of error. Such a probability, however, does not have a closed-form expression, but is upper bounded by the Bhattacharyya bound [Duda01-PatternClassification]. For a two-class GMM $\bm{X}\sim p_{0}\,\mathcal{N}(\bm{\mu}_{0},\,\bm{\Sigma}_{0})+p_{1}\,\mathcal{N}(\bm{\mu}_{1},\,\bm{\Sigma}_{1})$, the bound is
[…]
@@
where ||$\cdot$|| ⋅ | is the determinant of a matrix, and $\text{error}^{\star}$ is the classification error achieved by a Bayes classifier. We apply (25) to $\bm{X}$ and $\widehat{\bm{X}}$, whose models are in (12) and (13). Thus, $\bm{\mu}_{0}=\bm{0}_{n}$, $\bm{\mu}_{1}=\bm{D}\bm{E}\bm{c}_{n}$, and $\bm{\Sigma}_{0}=\bm{\Sigma}_{1}=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$. Hence,
$\mathbb{E}\Big{[}\epsilon_{c_{0}}(\bm{X},\,\widehat{\bm{X}})\Big{]}$ $=\mathbb{P}\big{(}\text{class}(\widehat{\bm{X}})\neq c_{0}\big{(}\widehat{\bm{X}}\big{)}\big{)}$
@@
So, in (18), rather than bounding $\mathbb{E}\big{[}\epsilon_{c_{0}}(\bm{X},\,\widehat{\bm{X}})\big{]}\leq C$, we impose instead that the right-hand side of (26) is upper bounded by $C$, which is equivalent to
@@
IV-B RDPCO: Heuristic algorithm for RDPC optimization
@@
Design of $\widehat{\bm{\Sigma}}$. While the original signals in (12) have non-degenerate distributions, the decoded signals in (13) have degenerate distributions. Specifically, assuming that $\bm{E}\in\mathbb{R}^{m\times n}$ and $\bm{D}\in\mathbb{R}^{n\times m}$ have full rank and that $\text{range}(\bm{E})\cap\text{null}(\bm{D})=\emptyset$, the output signals in (13) live in an $m$-dimensional subspace. If the fixed vector $\bm{c}_{n}$, which represents the distance between $\bm{X}|H_{0}$ and $\bm{X}|H_{1}$, is orthogonal to that subspace (equivalently $\bm{D}\bm{E}\bm{c}_{n}=\bm{0}_{n}$), then $\widehat{\bm{X}}|H_{0}$ and $\widehat{\bm{X}}|H_{1}$ become indistinguishable. In this case, classification is impossible and perception is also undermined [note that the second term in (24) requires $\|\bm{D}\bm{E}\bm{c}_{n}-\bm{c}_{n}\|_{2}$ to be small].
To avoid this, we first generate the (degenerate) covariance matrix $\widehat{\bm{\Sigma}}:=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$ by guaranteeing that the distance between $\bm{X}|H_{0}$ and $\bm{X}|H_{1}$ is preserved after transmitting these signals through the channel. We achieve this by guaranteeing that $\bm{c}_{n}$ is an eigenvector of $\widehat{\bm{\Sigma}}$ associated to eigenvalue $1$, while the remaining eigenvectors are associated to eigenvalues of smaller magnitude. Specifically, we set $\widehat{\bm{\Sigma}}=\bm{Q}\bm{\Lambda}\bm{Q}^{\top}$, where the first column of $\bm{Q}$ is $\bm{c}_{n}$ and the remaining ones are the output of Gram-Schmidt orthogonalization. Also, $\bm{\Lambda}=\text{Diag}(1,\lambda_{2},\ldots,\lambda_{m},0,\ldots,0)$, with $\lambda_{i}$ being drawn uniformly at random from $[0,1]$, for $i=2,\ldots,m$.
@@
Finding $(\bm{E},\bm{D})$. With $\widehat{\bm{\Sigma}}$ fixed and assuming that, at iteration $k$, $\bm{\Sigma}=\bm{\Sigma_{k-1}}$ is also fixed, we seek a factorization $\widehat{\bm{\Sigma}}=\bm{D}\bm{E}\bm{E}^{\top}\bm{D}^{\top}+\bm{D}\bm{\Sigma_{k-1}}\bm{D^{\top}}$. We do so via an intuitive process that leads to a unique factorization. Specifically, we design $\bm{E}$ and $\bm{D}$ such that $\bm{D}\bm{E}\bm{E}^{\top}\bm{D}^{\top}$ is as close to the identity matrix as possible (to preserve signals passing through the channel), while $\bm{D}\bm{\Sigma_{k-1}}\bm{D}^{\top}$ is as small as possible (to mitigate the effects of noise). Also, we ensure the principal direction $\bm{c}_{n}$ is preserved: $\bm{D}\bm{E}\bm{c}_{n}\simeq\bm{c_{n}}$. These requirements, weighted equally, can be cast as an optimization problem:
@@
[…]
@@
Finding $\bm{\Sigma}$. Once the encoder-decoder pair is fixed at $(\bm{E_{k}},\bm{D_{k}})$, we find the diagonal rate matrix $\bm{\Sigma}:=\text{Diag}(\bm{\sigma}):=\text{Diag}(\sigma_{1},\ldots,\sigma_{m})$ by applying a barrier method [Boyd04-ConvexOptimization] to (32), i.e., we solve a sequence of problems in $t$, each of which is
@@
where $\widehat{\bm{\Sigma}}_{\bm{k}}=\bm{D_{k}}\bm{E_{k}}\bm{E_{k}}^{\top}\bm{D_{k}}^{\top}+\bm{D_{k}}\bm{\Sigma}\bm{D_{k}}^{\top}$. In (40b), $D_{k}$ absorbs all the terms independent from $\bm{\Sigma}$ when we set $\mathbb{E}[\big{\|}\widehat{\bm{X}}-\bm{X}\big{\|}_{2}^{2}]\leq D$ in (22) [including $D$]. To obtain (40c), note that imposing the right-hand side of (24) to be smaller than $P$ is equivalent to $\big{\|}\widehat{\bm{\Sigma}}_{\bm{k}}^{\frac{1}{2}}-\bm{I_{n}}\big{\|}_{F}^{2}\leq\big{(}P-\|\bm{D_{k}}\bm{E_{k}}\bm{c_{n}}-\bm{c_{n}}\|_{2}\cdot p_{1}\big{)}^{2}=:P_{k}$. And $h_{p}(\bm{\Sigma})$ depends on $\bm{\Sigma}$ via $\widehat{\bm{\Sigma}}_{\bm{k}}$. Finally, (40d) is the direct application of the log-barrier function to (27).
[…]
@@
12:    Find $(\bm{E_{k}},\bm{D_{k}})$ via gradient descent applied to (37) [cf. (38)]